An Arctic tractor drags a 45 ton (that's metric ton 1 ton equals 1000 kg) across the snow which has a friction coefficient of 0.075 a distance of 4500 kilometers how much work did the tractor do
Answer by ikleyn(52878)

.

Convert all given data in kilograms and meters


    45 tons = 45000 kilograms (kg);

    4500 kilometers = 4500000 meters (m).


The friction force is the product of the weight by the friction coefficient k.


The weight is  W = 45000 kilograms * 10 m/s^2 = 450000 newtons.


The friction force is  F = k*W = 0.075*450000 = 33750 newtons.


The work done is the product of the friction force by the distance


    A = F*d = 33750 * 4500000 = 1.51875E+11 joules = 1.51875E+8  kilo-joules.   ANSWER

Solved.

--------------

What is your textbook in Physics ?